OpenStudy (anonymous):

What is an equation in point-slope form for (18, 26) and (20, 32)? What is the domain and range of the equation? Please explain the clearest and shortest way possible. Thanks!

OpenStudy (anonymous):

slope = 32-26 / 20-18 = 3 = m use y-y1 = m(x-x1) y-32 = 3(x-20) is that point-slope form? (we don't use that naming convention in the uk)

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

there you go then are you happy with that or do you need more explanaation